OpenStudy (anonymous):

how do i get a a/20+4/15=9/15

jhonyy9 (jhonyy9):

is the same with denominator common

jhonyy9 (jhonyy9):

how many will be for 15 and 20 ?

jhonyy9 (jhonyy9):

do you like to learn it ?

jhonyy9 (jhonyy9):

or you like getting ths answer ?

jhonyy9 (jhonyy9):

so ok

OpenStudy (anonymous):

the anwser i dont have time to learn it i have one more question

jhonyy9 (jhonyy9):

a/20=5/15

jhonyy9 (jhonyy9):

3a=20 a=20/3

jhonyy9 (jhonyy9):

because 60 is the denominator common

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

it give me 3/5

jhonyy9 (jhonyy9):

how ?

OpenStudy (anonymous):

cause i put 20/3 which =6.66666666666667 and i but the 20/3 where a is and it gave me 3/53..

OpenStudy (anonymous):

3/5

jhonyy9 (jhonyy9):

sorry but this style to solve it i not know and i think that is right how i have wrote to you

jhonyy9 (jhonyy9):

because when you have fractions with denominators different you need arrange those all to denominator common and just after can you make the needed works

jhonyy9 (jhonyy9):

assum or subtraction

OpenStudy (anonymous):

so whats A ? cause bro am confused

jhonyy9 (jhonyy9):

a=20/3

OpenStudy (anonymous):

can you put in the equation and it would give you 9/15??

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

20/3/20 =20/3 * 1/20 =1/3 yes ?

jhonyy9 (jhonyy9):

1/3 + 4/15 = (5+4)/15 =9/15

jhonyy9 (jhonyy9):

right ?

OpenStudy (anonymous):

so the anwser is 20/3?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

i have proven it to you

OpenStudy (anonymous):

thanks :)

jhonyy9 (jhonyy9):

not is understandably again to you ?

jhonyy9 (jhonyy9):

so if you think that i have helped to you in this exercise please push the good answer thank you